OpenStudy (anonymous):

Please help? Write an equation of a line that goes through point (4, –5) and is parallel to line y = 2x + 8. Show all work

OpenStudy (anonymous):

do you know what parallel lines have in common?

OpenStudy (anonymous):

yes the slope

OpenStudy (anonymous):

ok... so you know the line you want has to have the same slope as that given line... what's the slope of that given line? m= ???

OpenStudy (anonymous):

2 also right?

OpenStudy (anonymous):

yes... so now you have a point and slope... do you know point-slope form: \(\large y-y_1=m(x-x_1) \) ???

OpenStudy (anonymous):

just plug in those values.... \(\large (x_1, y_1) \rightarrow (4, -5) \) slope: \(\large m=2 \)

OpenStudy (anonymous):

Oh ok so 1 sec

OpenStudy (anonymous):

so y - (-5)= 2(x-4)+8?

OpenStudy (anonymous):

everything is correct except for that "+8" . where did that come from?

OpenStudy (anonymous):

Am i supposed to plug that in cause its in the other equation or no?

OpenStudy (anonymous):

no... just plug into this equation: \(\large y-y_1=m(x-x_1) \) this equation gives you the line with the same slope and passing through the given point. it's called point-slope form.

OpenStudy (anonymous):

Oh okay thanks!(:

OpenStudy (anonymous):

are you supposed to put it in slope-intercept form?

OpenStudy (anonymous):

yw.... :)

OpenStudy (anonymous):

all i had to do was simplify I got y=2x-13

OpenStudy (anonymous):

yep.... correct...:)